HUBBLE SPACE TELESCOPE DAILY REPORT #938 PERIOD COVERED: 0800 EST Daily Status Report as of 209/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: The following proposals were not executed due to the current Software Sunpoint Safemode event: (See Operations Note 2.4) 1.1 Completed FOS/BL 3441 (Post Asymptotic Giant Branch Evolution in the Magellanic Clouds) 1.2 Completed GHRS/2 5040 (GHRS Detector Calibrations - Cycle 3) 1.3 Completed Two Sets of WFC 4381 (HST Medium Deep Survey - Cycle 3) 1.4 Completed HSP/UV1 1091 (Ultraviolet Pulsations from X-Ray Pulsars) 1.5 Completed FOS/RD 5050 (Cycle 3 Internal Wavelenght Calibration) 1.6 Completed WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) 1.7 Completed Two Sets of PC 4929 (WF/PC Cycle 2 :Improve Dark Calibration) 1.8 Completed FOS/RD 4396 (QSO Absorption System Snapshot Survey - Part 1 of 3: Cycle 3 High) 1.9 Completed FOS/BL 4548 (Search for Evidence of Active Ice Volcanism of Europa Cycle 3 - High) 1.10 Completed FOS/RD/BL 4340 (The Origin of the Blue Featureless Continuum in Seyfert 2 Nuclei) 1.11 Completed PC 3912 (Cores of Early - Type Galaxies) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 0 Successful : 0 Scheduled Reacquisitions : 0 Successful : 0 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 8 Successful : 6 Updates of Trackers 1 & 2 failed the error box tests at 208/0710Z and 208/2143Z. Subsequent operations were okay. 2.3 Significant HSTARs: None 2.4 Operations Note: The HST entered software sunpoint safemode at approximately 207/0724Z ( Monday 26, 1993). Transition was successfully made to the Health and Safety load at 208/0600Z. The return to normal science observations is planned for 211/1200Z (Friday, July 30). Functional testing of the SADE/SA systems on-board the spacecraft was successfully completed this morning. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: The following proposals were scheduled to be executed, but will be canceled due to the current Software Sunpoint Safemode event: (See Operations Note 2.4) 3.1 Complete GHRS/2 5037 (GHRS Echelle Spectral Lamp Mini-Function Test) 3.2 Complete WFC 4809 (Cycle 2 : WF/PC PSF Calibration Proposal) 3.3 Complete PC 5021 (HST Observations of Comet Shoemaker-Levy (1993E)) 3.4 Complete FOS/RD 4581 (Quasar Absorption Line Survey - Cycle 3 High) 3.5 Uplink Updated Ephemeris, Per MOSES 2871 3.6 Complete WFC 4381 (HST Medium - Deep Survey Cycle 3) 3.7 Complete Two Sets of FOS/BL 4396 (QSO Absorption System Snapshot Survey - Part 1 of 3: Cycle 3 High) 3.8 Complete Two Sets of WFC 5013 (WF/PC Cycle 2 Calibration: WFC Flats Narrow. 2) 3.9 Complete GHRS/2 5036 (GHRS 1st Order Grating Spectral Lamp Test - Cycle 3) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: The Type 10 (EVA 4) Simulation will be conducted tomorrow (Thursday, July 29) from 11:00 A.M. until Midnight (EDT). /DMH -END-